OpenStudy (anonymous):

is cos inverse (cos^-1) the same thing as secant?

OpenStudy (anonymous):

No. Inverse trig functions return angles. secant will return a length as will all trig functions.

OpenStudy (anonymous):

I made this sheet a while ago, i think you will find this fairly useful for trig. http://learnix.net/ultimate-trig-cheat-sheet/

OpenStudy (anonymous):

It's easy to confuse arccos, AKA inverse cosine with the algebraic notation of x^(-1). In the case of trig functions, it means something else...it means that it's the inverse function. Does that make sense?

OpenStudy (anonymous):

ok thanks! then how do you solve the problem \[\cos^{-1} (\cos (17\pi/5))\]

OpenStudy (anonymous):

@agentc0re

OpenStudy (anonymous):

First you would need to calculate the inner part.

OpenStudy (anonymous):

@agentc0re ok but how do i calculate it without a calculator since it is not a unit circle value

OpenStudy (anonymous):

wait isnt \[\cos^{-1} \cos \] just one?

OpenStudy (anonymous):

Well, not 1 but they do kind of "cancel" each other out.

OpenStudy (anonymous):

ok thank you so much!! so then the answer would just be -17pi/5

OpenStudy (anonymous):

oops not the negative

OpenStudy (anonymous):

Here's the problem though. You're going to get an answer that is outside of the unit circle. If you try to convert that angle to a degree it's going to be much larger than 360.

OpenStudy (anonymous):

So you'll need to be able to devise a way to know where you expect it to be on the unit circle. Make sense?

OpenStudy (anonymous):

YES THANK YOU!!!!!!!!!!!!!!!

OpenStudy (anonymous):

Awesome! Good luck :D